OpenStudy (anonymous):

What is the value of cos(5π/3)?

OpenStudy (anonymous):

\[ \cos(5\pi/3) = \cos(5\pi/3 - 2\pi) = \cos(5\pi/3-6\pi/3) = \cos(-\pi/3) = \cos(\pi/3) \]

OpenStudy (anonymous):

\[\pi/3 \cdot 180^\circ/\pi = 180^\circ/3 = 60^\circ \]

OpenStudy (anonymous):

Than you! @wio

OpenStudy (anonymous):

Thank*

OpenStudy (anonymous):

Well, I didn't give you the answer... do you know what \(\cos 60^\circ\) is?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

It is supposed to be equilateral but I didn't draw perfect

OpenStudy (anonymous):

it equals 1/2

OpenStudy (anonymous):

Thank you for your help!